Mathslinger:

Recall that the perimeter of a rectangle is P=2(W+L), where W is the width and L is the length. The length of a rectangle is 24 feet more than the width. If the perimeter is 80 feet, then what is the length of the rectangle, in feet?

surjithayer:

l=w+24 p=2(l+w)=2(w+24+w)=80 2w+24=80/2=40 find w and then l